Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
rolland - 18:09 Monday 20 June 2022 (56199) Print this report
NE and WE PCal laser switched off

The laser of the two PCal have been switched off on purpose at 16h UTC, to allow the activity of tomorrow morning on the PCal benches (modification of the cover of the NE injection bench to have less risk to touch the NCal when opening/closing the PCal bench ; tests of a cap design around the viewports to better stop scattered light between PCal benches and end towers). The plot shows the offsets of the four photodiode channels (few minutes after the stop).

Images attached to this report
Comments to this report:
lieunard, journet, rolland - 18:19 Wednesday 22 June 2022 (56228) Print this report

The planned activities have been done yesterday during the maintenance:

  • tests of design for PCal viewport covers. One piece has been tested on the four viewports (WE Tx, WRx, NE Tx, NE Rx). Some tunings of the design will be done to prepare all the covers for a later installation. The diaphragms that were in front of the viewports have been removed for this test. They will be brought back to LAPP for integration in the new covers.
  • modification of the cover of the NE injection bench, close to the NCal (see pictures 1 to 4): the cover has been cut into two pieces horizontally to allow opening/closing the bench without touching the neighbouring NCal support and rotors. This activity has required to access some machining tools in the building 1500W. The planned tool was faulty but a fall-back solution has been found to machine the cover. It was put back in place on NE bench today around 14h-14h30 UTC.

The PCal lasers have been switched on today:

  • WE Pcal around 14h50 UTC, power set at 1.3 W, in loop on Tx_PD2, filter1. See figures 5 (laser off), 6 (laser on, open loop) and 7 (closed loop).
  • NE Pcal around 16h UTC, power set at 1.3 W, in loop Tx_PD2, filter1. See figures 8(laser off), 9 (laser on, open loop) and 10 (closed loop). The cross-calibration of PD1 and PD2 has changed a lot. PD1 power is much less than PD2, which indicates that at least PD1 has been misaligned during the intervention. Also, looking at the noise level in closed loop, there are bumps similar (even in larger) to bumps seen when the beam was clipped towards a photodiode.  -> The NE Pcal bench need to be recalibrated.  To check if a rough calibration of PD2 is still fine,it will be compared with the laser command before and after the intervention.

 

 

Images attached to this comment
Search Help
×

Warning

×